# Do Large Language Model Understand Multi-Intent Spoken Language ?

### v2

@@ Abstract
This study marks a significant advancement by harnessing Large Language Models (LLMs) for multi-intent spoken language understanding (SLU), proposing a unique methodology that capitalizes on the generative power of LLMs within an SLU context. Our innovative technique reconfigures entity slots specifically for LLM application in multi-intent SLU environments and introduces the concept of Sub-Intent Instruction (SII), enhancing the dissection and interpretation of intricate, multi-intent communication within varied domains. The resultant datasets, dubbed LM-MixATIS and LM-MixSNIPS, are crafted from pre-existing benchmarks. Our research illustrates that LLMs can match and potentially excel beyond the capabilities of current state-of-the-art multi-intent SLU models. It further explores LLM efficacy across various intent configurations and dataset proportions. Moreover, we introduce two pioneering metrics, Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA), to provide an in-depth analysis of LLM proficiency in this complex field.
@@ 1 Introduction
Spoken language understanding (SLU) is a crucial component of task-oriented dialog systems, which aim to create a semantic frame that summarizes the user’s request. The semantic frame is specially constructed using intent detection to identify user’s intents and slot filling to extract relevant semantic constituents. Since the two sub-tasks of intent detection and slot filling are closely tied (Tur and Mori, 2011), dominant SLU systems adopt joint models to model the correlation between them (Liu and Lane, 2016; Goo et al., 2018; Qin et al., 2019).
@@
Figure 1: An example with multi-intent detection and slot filling and “TP” denote “Transition Point”.
To deal with multi-intent scenarios, an increasing number of studies have begun to focus on modeling SLU in multi-intent settings. Xu and Sarikaya (2013) and Kim et al. (2017) first explored the multi-intent SLU. Then Qin et al. (2020a) proposed an adaptive interaction framework (AGIF) to achieve fine-grained multi-intent information integration for token-level slot filling. Qin et al. (2021b) further proposed a global-locally graph interaction network (GL-GIN) to model slot dependency and interaction between multiple intents. Recently, Huang et al. (2022) proposed a chunk-level intent detection (CLID) framework to split multi-intent into single-intent with an intent transition point, achieving a promising performance.
@@
Addressing this, we pose several critical research questions: RQ1: Can Large Language Models effectively handle multi-intent SLU, and if so, how can they be adapted or constructed to fit within a multi-intent SLU framework? RQ2: How does the performance of Large Language Models compare with the current state-of-the-art in multi-intent SLU? RQ3: Does the size of a Large Language Model influence its performance in multi-intent SLU? RQ4: What methods can be employed to fully evaluate the performance of Large Language Models in this task?
In pursuit of answers to these questions, we propose a novel methodological approach that harnesses the strengths of LLMs for multi-intent SLU. This involves the adaptation of entity slots from conventional models to accommodate a unified framework designed for simultaneous multiple intent detection and slot filling. This approach seeks to fully harness the potential of LLMs in multi-intent task performance. Moreover, drawing from the pivotal influence of sub-sentence elements Huang et al. (2022), we introduce the Sub Intent Instruction (SII) concept. This methodology enhances LLMs’ capacity for discerning and partitioning multi-intent utterances into individual sub-intent segments, dovetailing with ’chain of thought’ reasoning Wei et al. (2022).
@@
For detailed information on related work, please refer to Appendix A. Our code and datasets are available at https://github.com/SJY8460/SLM.
@@ 2.2 Slot Filling
Figure 2: Diagram illustrating the workflow of LLM training and inference. "Optional" denotes that sub-intent instructions are utilized only in additional explorations, not the main experiment.
@@ 3 Approach
As depicted in Figure 2, our methodology establishes a detailed framework for the integration of LLMs within the sphere of multi-intent SLU. This framework is composed of several essential components: the innovative reconfiguration of conventional slot labels to create entity slots; the introduction of Sub Intent Instruction, which bolsters the model’s grasp of complex intents; a dedicated training stage featuring supervised fine-tuning to refine model performance; the incorporation of QLoRA Quantization to streamline model scaling; and a comprehensive inference procedure to guarantee precise multi-intent SLU discernment.
@@ 3.1 Entity Slots Construction
The process of transforming traditional BIO-annotated slots into an entity-slot mapping is mathematically articulated as follows: Consider a sequence of tokens $T=\{t_{1},t_{2},...,t_{n}\}$ alongside a corresponding sequence of BIO-annotated slots $S=\{s_{1},s_{2},...,s_{n}\}$. The transformation into entity-slot mapping is encapsulated by a function $f$:
$f(T,S)\mapsto E$ (1)
@@
where $E=\{e_{1},e_{2},...,e_{m}\}$ represents the resultant entity-slot mapping and $m$ donates the number of entity slot within an utterance. The function $f$ processes the sequence $S$, identifying ’B-’ as the initiation of an entity and ’I-’ as its continuation. The transition to an ’O’ tag or a shift in the entity type signals the conclusion of the current entity. At this juncture, the accumulated tokens corresponding to this entity are amalgamated and incorporated into $E$ as a key-value pair. The key represents the entity type, while the value constitutes the concatenated tokens. This procedure iterates until all entities within $S$ are systematically extracted and mapped.
3.2 Sub Intent Instruction
Inspired by Huang et al. (2022)’s method of chunk-level intent detection (CLID), which segments multi-intent utterances into individual intents, we refine our model’s comprehension of intents and slots within sentences. Our approach, grounded in the practice of deconstructing multi-intent utterances into singular intents, introduces the concept of ’sub-intent’ guidance during training. By employing a chain-of-thought concept (Wei et al., 2022), our objective is to enhance the model’s capability to accurately identify and classify intents and slots.
In particular, our training dataset is enriched with sub-utterance instructions that prompt the model to concentrate on the intricate intent and slot configurations present in multi-intent utterances. This heightened level of detail during training fosters a more nuanced perception and response formulation. The training samples encompass not only the primary utterance but also a clear directive for sub-utterances associated with each respective intent.
As depicted in Figure 2, our instructional strategy mirrors the human cognitive approach of parsing complex inquiries into more manageable segments, thereby directing the model to emulate a corresponding thought process. By incorporating these sub-utterance directives, we aim to scrutinize their influence on improving LLMs’ efficacy in SLU tasks.
3.3 Supervised Fine-tuning
We utilize supervised fine-tuning to refine the generative capabilities of LLMs, ensuring they align with the detailed structures of multi-intent SLU. Initially, LLMs are trained on a broad range of data, which imparts a general understanding of language. However, to apply these models effectively in the specialized domain, supervised fine-tuning is essential. It involves the iterative adjustment of the pre-trained model’s parameters using a task-specific labeled dataset. This targeted refinement enhances the model’s accuracy and its ability to recognize and replicate the nuanced patterns unique to SLU tasks.
The fine-tuning process is defined mathematically as an optimization problem. The goal is to identify the optimal parameter set $\theta^{*}$ that minimizes the loss function $L$ over the training dataset $D$. This is expressed in the following equation:
[…]
In this context, the loss function $L$ is typically instantiated as a cross-entropy loss, which quantifies the discrepancy between the model’s predictions and the actual labels in the dataset $D$. This process is facilitated by back-propagation, which computes gradients that guide the adjustment of parameters $\theta$ towards minimizing the loss. Through this meticulous process of supervised fine-tuning, LLMs are transformed from their generalist pre-training state to become adept at handling the specific demands of multi-intent SLU.
3.4 Efficient Fine-tuning with QLoRA Quantization
In optimizing LLMs for multi-intent SLU, our approach integrates Quantized Low-Rank Adapters (QLoRA) Dettmers et al. (2023), markedly improving both efficiency and scalability. QLoRA, an advancement on the Low-Rank Adapters (LoRA) technique Hu et al. (2022a), selectively modifies a limited set of model weights. This targeted adaptation significantly reduces memory usage during training, enhancing resource management. Further, QLoRA employs quantized weights in the LoRA fine-tuning process, which substantially lowers the memory footprint.
This integration of QLoRA into our LLM fine-tuning strategy leads to a process that is not only memory-efficient but also maintains high performance, crucial for the nuanced demands of multi-intent SLU. By adopting this approach, we harness advanced capabilities of LLMs even in environments with limited computational resources, making sophisticated NLP technologies more accessible and scalable.
3.5 Inference Phase
@@
During the inference phase for our LLMs in multi-intent SLU, we focus on processing test sentences to extract intents and entity slots. This process involves utilizing regular expressions to parse the LLM’s output, followed by a structured transformation into the standard BIO tagging format, as shown in Figure 2. The transformation process is mathematically formulated to ensure precision and clarity. Given a sentence represented by a sequence of words $T=\{t_{1},t_{2},...,t_{n}\}$, we initially tag each word with ’O’, denoting the absence of an entity. Entity slots extracted from the LLM output are represented as $E=\{e_{1},e_{2},...,e_{m}\}$. The task then is to align these slots with the words in $T$ and reassign appropriate BIO tags. The re-tagging function, denoted as $f$, operates on $T$ and $E$ to produce a sequence of BIO tags $S=\{s_{1},s_{2},...,s_{n}\}$. The function is defined as:
$f(T,E)\rightarrow S$ (3)
@@
SDJN(BERT) 87.5 78.0 46.3 95.4 96.7 79.3
CLID(Roberta) 85.9 80.5 49.4 96.0 97.0 82.2
ChatGPT (5-shot) 64.0 54.1 14.6 62.9 83.9 12.7
Vicuna-7B-v1.5 Peng et al. (2023) 83.3 79.5 47.3 95.7 97.6 78.9
Llama-2-7B-chat Touvron et al. (2023) 86.5 82.4 51.1 95.7 96.9 78.9
Vicuna-13B-v1.5 Peng et al. (2023) 87.9 $\textbf{83.6}^{*}$ 50.8 95.9 97.5 80.7
Llama-2-13B Touvron et al. (2023) 87.9 81.0 49.5 96.7 97.8 $\textbf{83.3}^{*}$
Mistral-7B-Instruct-v0.1 Jiang et al. (2023) 88.7 80.6 $\textbf{53.4}^{*}$ 95.6 97.6 79.8
Table 1: SLU performance on MixATIS and MixSNIPS datasets. Values with * indicate that the improvement from our model is statistically significant over all baselines ($p<0.05$ under t-test).
Model intent num = 1 intent num = 2 intent num = 3
CLID 88.6 94.7 76.4 88.1 77.5 48.4 87.6 64.3 28.5
CLID(Roberta) 88.6 95.8 77.6 85.4 80.3 48.8 84.7 66.8 29.0
Vicuna-7B-v1.5 90.0 94.4 80.4 86.7 82.7 50.9 85.2 70.5 28.0
Llama-2-7B-chat 86.9 96.5 79.7 87.0 83.1 50.9 85.6 70.5 31.0
Mistral-7B-Instruct-v0.1 90.0 93.0 77.6 88.9 81.6 55.3 88.0 69.0 31.5
@@ 4.3 Baselines
In our study, we benchmark LLMs performance against a range of established baselines in the multi-intent SLU domain. These include vanilla models like Stack-Propagation (Qin et al., 2019): a stack-propagation framework to explicitly incorporate intent detection for guiding slot filling. AGIF (Qin et al., 2020b): an adaptive interaction network to achieve fine-grained multi-intent information integration, GL-GIN (Qin et al., 2021b): a local slot-aware and global intent-slot interaction graph framework to model the interaction between multiple intents and all slots within an utterance, SDJN (Chen et al., 2022): a multiple instance learning and self-distillation framework for weakly supervised multiple intent information capturing, CLID (Huang et al., 2022): a chunk-level intent detection framework for recognizing intent within a fragment of an utterance and SSRAN (Cheng et al., 2023): a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection. We also included PLM-based models like SDJN(Bert) and CLID(Roberta) which are extensions of their respective base models with pre-train language model backbones.
@@ 4.4 Large Language Model
The landscape of natural language processing has been revolutionized by the advent of open-source Large Language Models such as Llama2(Touvron et al., 2023), VicunaPeng et al. (2023), and MistralJiang et al. (2023). These models’ advanced parameterization and diverse training regimes have notably enhanced NLP task performance.
@@
In our study, we harness state-of-the-art LLMs from these open-source initiatives, including configurations such as Vicuna-7B-v1.5 and Vicuna-13B-v1.5 (Peng et al., 2023), Llama-2-7B-chat and Llama-2-13B, and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023). For comprehensive details on these LLMs, please consult Appendix A.2.
Moreover, we extend our evaluation to examine ChatGPT’s proficiency in a 5-shot learning context, paralleling the tests conducted with the LLMs outlined above. A distinctive aspect of our methodology with ChatGPT involves introducing an explicit intent and slot list to direct the model’s learning, enhancing its capability to align with specified NLP tasks.
4.5 Main Results Analysis
@@
We evaluate performance in these areas: F1 score is used for slot filling, accuracy for intent prediction, and overall accuracy for sentence-level semantic frame parsing according to Qin et al. (2021b) and Huang et al. (2022). The overall accuracy represents the proportion of sentences in which both intent and slot are accurately predicted and it is the most important metric in multi-intent SLU.
As detailed in Table 1, our results indicate that: (1) The LLMs have manifested notable enhancements when compared with the established model CLID(Roberta). Within the single-domain MixATIS dataset, Mistral-7B-Instruct-v0.1 outstrips CLID(Roberta) by exhibiting a 2.8% increase in intent prediction accuracy, a 0.1% rise in slot filling F1 score, and a 4.0% advancement in overall sentence-level semantic frame parsing accuracy. In the multi-domain MixSNIPS dataset, Llama2-13B shows improvement over CLID(Roberta) with a 0.7% boost in intent prediction accuracy, a 0.8% increase in slot filling F1 score, and a 1.1% growth in overall accuracy. (2) When comparing the performance based on the size of the LLMs, the larger 13B models do not uniformly outperform the 7B models in the MixATIS dataset. This suggests that the model size alone may not resolve the complexity of detailed slot and intent recognition in single-domain datasets. However, the scenario is different in the multi-domain MixSNIPS dataset, where the expanded capabilities of the 13B models seem to confer an advantage. Specifically, Vicuna-13B-v1.5 improves upon its 7B counterpart by 1.8% in overall accuracy, and the Llama2-13B notably enhances performance by 4.4% in the same metric. This observation suggests that larger models can be more effective in environments with a wide range of domains, likely due to their superior ability to capture and process the diverse and intricate patterns present in the data.

### v3

@@ Abstract
This research signifies a considerable breakthrough in leveraging Large Language Models (LLMs) for multi-intent spoken language understanding (SLU). Our approach re-imagines the use of entity slots in multi-intent SLU applications, making the most of the generative potential of LLMs within the SLU landscape, leading to the development of the EN-LLM series. Furthermore, we introduce the concept of Sub-Intent Instruction (SII) to amplify the analysis and interpretation of complex, multi-intent communications, which further supports the creation of the ENSI-LLM models series. Our novel datasets, identified as LM-MixATIS and LM-MixSNIPS, are synthesized from existing benchmarks. The study evidences that LLMs may match or even surpass the performance of the current best multi-intent SLU models. We also scrutinize the performance of LLMs across a spectrum of intent configurations and dataset distributions. On top of this, we present two revolutionary metrics - Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA) - to facilitate a detailed assessment of LLM competence in this multifaceted field." Our code and datasets are available at https://github.com/SJY8460/SLM.
@@ 1 Introduction
Spoken language understanding (SLU) is a crucial component of task-oriented dialog systems, which aim to create a semantic frame that summarizes the user’s request. The semantic frame is specially constructed using intent detection to identify user’s intents and slot filling to extract relevant semantic constituents. Since the two sub-tasks of intent detection and slot filling are closely tied (Tur and Mori, 2011), dominant SLU systems adopt joint models to model the correlation between them (Goo et al., 2018; Qin et al., 2019).
Figure 1: An example with Multi-Intent SLU, where B-WT donates B-Weather, B-LOC donates B-Location and “TP” denote “Transition Point”.
@@
Recently, Huang et al. (2022) proposed a chunk-level intent detection (CLID) framework to split multi-intent into single-intent with an intent transition point. Furthermore, Yin et al. (2024) develop an united multi-view intent-slot interaction framework(Uni-MIS), achieving promising performance.
@@
Addressing this, we pose several critical research questions: RQ1: Can Large Language Models effectively handle multi-intent SLU, and if so, how can they be adapted or constructed to fit within a multi-intent SLU framework? RQ2: How does the performance of Large Language Models compare with the current state-of-the-art in multi-intent SLU? RQ3: What methods can be employed to fully evaluate the performance of Large Language Models in this task?
To address these queries, we introduce a novel method, leveraging the capabilities of LLMs for multi-intent SLU. This method adapts the entity slots from traditional models, thus offering a unified framework designed for simultaneous multiple intent detection and slot filling. As depicted in Fig 1, the task of token-level slot labeling, often termed as slot filling, can be further refined into an entity-level slot detection. This enables LLMs to focus on identifying pertinent slot labels without the necessity to label all slot tokens within a sentence, thereby mitigating the problem of uncontrolled length generation. Furthermore, driven by "Chain of Thought" reasoning Wei et al. (2022), we usher in the concept of Sub Intent Instruction (SII). This strategy enhances LLMs’ proficiency in distinguishing and segregating multi-intent utterances into individual sub-intent segments. For a more comprehensive evaluation of our proposed method, we introduce two innovative metrics: Entity Slot Accuracy (ESA) and Combined Semantic Accuracy (CSA). These metrics offer a thorough and insightful assessment of LLMs in the context of multi-intent SLU at the entity-level.
@@ 2.2 Slot Filling
Figure 2: Diagram illustrating the workflow of LLM training and inference. Sub-Intent instruction is only used in ENSI models.
@@ 3 Approach
As depicted in Figure 2, our methodology establishes a detailed framework for the integration of LLMs within the sphere of multi-intent SLU. This framework is composed of several essential components: the innovative reconfiguration of conventional slot labels to create entity slots; the introduction of Sub Intent Instruction, which bolsters the model’s grasp of complex intents; and a comprehensive inference procedure to guarantee precise multi-intent SLU discernment.
@@ 3.1 Entity Slots Construction
The process of transforming traditional BIO-annotated slots into an entity-slot mapping is mathematically articulated as follows: Consider a sequence of tokens $T=\{t_{1},t_{2},...,t_{n}\}$ alongside a corresponding sequence of BIO-annotated slots $S=\{s_{1},s_{2},...,s_{n}\}$. The transformation into entity-slot mapping is encapsulated by a function $c$:
$c(T,S)\mapsto E$ (1)
@@
where $E=\{e_{1},e_{2},...,e_{m}\}$ represents the resultant entity-slot mapping and $m$ donates the number of entity slot within an utterance. The function $c$ processes the sequence $S$, identifying ’B-’ as the initiation of an entity and ’I-’ as its continuation. The transition to an ’O’ tag or a shift in the entity type signals the conclusion of the current entity. At this juncture, the accumulated tokens corresponding to this entity are amalgamated and incorporated into $E$ as a key-value pair. The key represents the entity type, while the value constitutes the concatenated tokens. This procedure iterates until all entities within $S$ are systematically extracted and mapped.
3.2 Sub-Intent Instruction
Inspired by the process of chunk-level intent detection (CLID)(Huang et al., 2022), which breaks down multi-intent utterances into singular intentions, and guided by the "Chain of Thought" reasoning Wei et al. (2022), we refine our model’s discernment of intents within sentences. Our methodology is grounded in the systematic dissection of multi-intent statements into distinct intents. For an instance dealing with two intents, the Sub-Intent Instruction could be represented as {"Sub-Intent-1": Sub-Utterance-1; "Sub-Intent-2": Sub-Utterance-2}. This enhanced granularity during the training phase fosters a more refined understanding and formulation of responses. Hence, the training samples not only include the primary utterance but also precise directives for sub-utterances correlated with each individual intent, facilitating the opportunity for improved comprehension and professionalism..
3.3 Efficient Fine-tuning with LoRA
In an effort to generate more cost-efficient experiments, we leverage advancements such as Low-Rank Adapters (LoRA) Hu et al. (2022a). Our attention shifts towards constraining the modifications in an existing weight matrix $W_{0}\in R^{d\times k}$. Through a low-rank decomposition strategy, these alterations are maneuvered towards $W_{0}+\Delta W=W_{0}+BA$, where $B\in R^{d\times r}$,$A\in R^{r\times k}$, and and the rank $r\ll min(d,k)$. As the training process progresses, the consistency of $W_{0}$ is maintained, with the flexible parameters being assigned to matrices $A$ and $B$ exclusively. This methodology culminates in an adapted forward pass, defined as:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (2)
Maintaining true to LoRA’s approach, the initialization stage employs a random Gaussian for A and zero for B, leading to $\Delta W=BA$ starting from zero at inception. The selective modification of a restricted set of model weights keeps the core structure intact, whilst providing ample room for focused enhancements. To further conserve computational resources for larger scale models, we also incorporate quantization techniques (Dettmers et al., 2023).
3.4 Extract and Recovery Inference
During the inference phase, LLMs in multi-intent SLU are primarily engaged in processing test sentences for the extraction of intents and entity slots. Notably, this process leverages the utilization of regular expressions to parse the first appearing intent and entity slot information from the LLM’s output, which is necessary due to the unpredictable generation length of the LLMs. This information is subsequently obligingly transformed in a structured manner into the standard BIO tagging format, a process illustrated in Figure 2.
@@ 3.4 Extract and Recovery Inference
The transformation process is mathematically formulated to ensure precision and clarity. Given a sentence represented by a sequence of words $T=\{t_{1},t_{2},...,t_{n}\}$, we initially tag each word with ’O’, denoting the absence of an entity. Entity slots extracted from the LLM output are represented as $E=\{e_{1},e_{2},...,e_{m}\}$. The task then is to align these slots with the words in $T$ and reassign appropriate BIO tags. The re-tagging function, denoted as $r$, operates on $T$ and $E$ to produce a sequence of BIO tags $S=\{s_{1},s_{2},...,s_{n}\}$. The function is defined as:
$r(T,E)\rightarrow S$ (3)
@@
SDJN${}_{\text{(Bert)}}$ (Chen et al., 2022) 87.5 78.0 46.3 95.4 96.7 79.3
CLID${}_{\text{(Roberta)}}$ (Huang et al., 2022) 85.9 80.5 49.4 96.0 97.0 82.2
Uni-MIS${}_{\text{(Roberta)}}$ Yin et al. (2024) 88.3 78.5 52.5 96.4 97.2 83.4
EN-Vicuna (Ours) 83.3 79.5 47.3 95.7 97.6 78.9
EN-Llama-2 (Ours) 86.5 82.4 51.1 95.7 96.9 78.9
EN-Mistral (Ours) 88.7 80.6 $\textbf{53.4}^{*}$ 95.6 97.6 79.8
ENSI-Vicuna (Ours) 86.0 80.6 49.4 96.2 97.1 82.4
ENSI-Llama-2 (Ours) 84.7 79.3 46.6 96.5 97.6 83.5
ENSI-Mistral (Ours) 86.4 78.0 49.8 96.7 97.5 83.9
Table 1: Multi-Intent SLU performance on MixATIS and MixSNIPS datasets. Values with * indicate that the improvement from our model is statistically significant over all baselines ($p<0.05$ under t-test).
@@ 4.3 Baselines
In our study, we benchmark LLMs performance against a range of established baselines in the multi-intent SLU domain. These include vanilla models like Stack-Propagation (Qin et al., 2019): a stack-propagation framework to explicitly incorporate intent detection for guiding slot filling. AGIF (Qin et al., 2020b): an adaptive interaction network to achieve fine-grained multi-intent information integration, GL-GIN (Qin et al., 2021b): a local slot-aware and global intent-slot interaction graph framework to model the interaction between multiple intents and all slots within an utterance, SDJN (Chen et al., 2022): a multiple instance learning and self-distillation framework for weakly supervised multiple intent information capturing, CLID (Huang et al., 2022): a chunk-level intent detection framework for recognizing intent within a fragment of an utterance and SSRAN (Cheng et al., 2023): a transformative network built on the Transformer model, designed to reduce the complexity of multi-intent detection in SLU through scope recognition and bidirectional interaction between results of slot filling and intent detection. We also included PLM-based models such as Uni-MIS (Yin et al., 2024): a unified multi-intent slu framework via multi-view intent-slot interaction. Additionally, SDJN(Bert) and CLID(Roberta) extend their respective base models by incorporating pre-trained language model backbones.
@@ 4.4 Large Language Model
In our study, we harness state-of-the-art LLMs from these open-source initiatives, including configurations such as Vicuna-7B-v1.5(Peng et al., 2023), Llama-2-7B-chat and Mistral-7B-Instruct-v0.1 (Jiang et al., 2023). We also report the performance of larger scale open source language models and ChatGPT(few-shot) in appendix A.
4.5 Detailed Analysis of Main Results
@@
For a robust evaluation of performance criteria, various metrics were adopted, namely: the F1 score for slot filling, accuracy for intent prediction, and overall accuracy for sentence-level semantic frame parsing, in alignment with the methodologies proposed by Qin et al. (2021b) and Huang et al. (2022). The paramount metric in multi-intent SLU is the overall accuracy, representing the proportion of sentences where both intent and slot are precisely predicted.
In this study, the term ‘EN-LLM model’ refers to LLM adapted with an entity slot, whereas ’ENSI-LLM’ models incorporate additional instructions for sub-intent. As manifested in Table 1, our findings underscore several observations:
(1) LLMs have demonstrated significant advancements compared to the prevalent model Uni-MIS. For instance, in the single-domain MixATIS dataset, EN-Mistral surpasses Uni-MIS by demonstrating a 0.4% enhancement in intent prediction accuracy, accompanied by a 2.1% improvement in slot filling F1 score, and a 0.9% uplift in overall sentence-level semantic frame parsing accuracy. Similarly, within the multi-domain MixSNIPS dataset, ENSI-Mistral outperforms Uni-MIS, reflecting a 1.1% increase in intent prediction accuracy and a 0.5% surge in overall accuracy. Our LLM models’ performance has also proven impressive when benchmarked against other formidable baseline models.
(2) With the integration of SII in the MixSNIPS datasets, significant growth is observed in overall accuracy rates. ENSI-Vicuna’s Overall (Acc) records a substantial increase of 3.5%, elevating from 78.9% to 82.4%. ENSI-Llama-2 follows a similar trajectory with a robust enhancement of 4.6%, achieving an Overall(Acc) of 83.5%. ENSI-Mistral benefits from a critical gain of 4.1% in Overall(Acc), culminating at 83.9%. However, a contrast is seen within the MixATIS dataset, where a slight drop is observed. This suggests that while SII proves beneficial for multi-domain tasks, it is less suited for single-domain tasks due to knowledge complexity. An LLM, when faced with an array of similar information, can lose accuracy in the prediction of intents and slots.
Model intent num = 1 intent num = 2 intent num = 3
CLID(Roberta) 88.6 95.8 77.6 85.4 80.3 48.8 84.7 66.8 29.0
Uni-MIS 89.2 95.1 78.6 87.6 78.3 50.5 86.7 66.7 31.7
EN-Vicuna 90.0 94.4 80.4 86.7 82.7 50.9 85.2 70.5 28.0
EN-Llama-2 86.9 96.5 79.7 87.0 83.1 50.9 85.6 70.5 31.0
EN-Mistral 90.0 93.0 77.6 88.9 81.6 55.3 88.0 69.0 31.5
Model intent num = 1 intent num = 2 intent num = 3
EN-Vicuna 95.1 97.6 88.4 96.0 97.9 79.9 95.4 96.8 67.6
EN-Llama 95.5 97.3 89.3 96.0 97.1 80.6 95.0 95.8 65.0
EN-Mistral 95.8 97.8 90.2 96.1 98.1 81.3 94.8 96.2 65.8
Table 3: The result comes from the dataset MixSNIPS. The intent num denotes the number of intents in an utterance.